# LLMEval-3: A Large-Scale Longitudinal Study on Robust and Fair Evaluation of Large Language Models

### v3

@@ Appendix C LLMEval-3 Leaderboard
Qwen-3-235B 96.43 90.32 9.23 9.43 9.03 9.50 8.23 9.43 8.97 9.17 8.73 8.60
QwQ-32B 94.49 88.54 8.30 9.46 9.23 9.33 7.83 9.46 8.65 9.27 8.57 8.43
GLM-4-32B 88.42 82.83 7.77 8.97 8.33 8.33 7.03 9.13 8.27 8.77 8.23 8.00
Qwen2.5-32B-Instruct 85.04 79.68 7.70 8.57 8.33 8.33 6.70 8.50 8.17 7.70 7.60 8.08
Qwen-Turbo-1101 83.68 78.41 7.97 8.37 8.03 8.23 6.40 8.50 8.10 7.50 7.27 8.05
Yi-34B-Chat 70.14 65.70 5.77 6.63 7.37 7.53 5.47 5.77 5.47 7.47 6.30 7.93
Nanbeige-plus 65.12 61.00 5.78 5.57 6.77 7.37 5.37 5.93 5.45 6.30 5.67 6.77
Baichuan2-13B-Chat 58.29 54.60 4.47 5.53 7.40 6.90 4.63 4.80 4.33 6.23 4.60 5.70
Llama-3.3-70B 60.84 57.00 5.80 6.90 5.63 5.70 5.47 5.70 6.30 4.70 4.87 5.93
Llama-3.2-90B-Vision 61.76 57.83 5.63 6.33 6.20 5.80 4.73 6.10 6.57 5.03 5.27 6.17
Qwen-plus 56.59 53.00 4.40 5.10 6.53 6.53 5.00 4.77 4.87 5.17 5.13 5.50
@@
Nanbeige-16B 55.44 51.93 4.37 5.30 6.50 6.30 3.97 4.70 4.07 5.90 4.73 6.10
Mixtral–8x7B-Instruct 51.67 48.40 4.27 5.47 6.47 6.40 3.13 4.50 5.07 3.57 4.37 5.17
ChatGLM2-6B 42.31 39.63 2.33 3.77 5.97 6.13 2.83 3.83 2.60 3.80 4.00 4.37
ziya_v1.1-13b 40.17 37.63 2.77 3.97 5.17 5.33 2.80 3.77 2.53 3.70 3.03 4.57
InternLM-Chat-7B 38.72 36.27 2.63 3.67 4.87 5.57 3.17 3.33 2.33 4.03 3.13 3.53
Phi-3-Medium-128K 36.94 34.60 2.27 4.17 3.70 4.23 2.87 4.50 3.57 3.20 2.27 3.83
Linly-Chinese-LLaMA-2-13B-hf 37.05 34.70 2.20 3.77 4.50 5.00 2.43 3.33 2.53 3.90 2.50 4.53
BELLE-Llama2-13B-chat-0.4M 36.26 33.97 2.57 3.07 4.93 4.73 2.83 3.80 2.43 3.33 2.40 3.87
Llama-2-7b-chat-hf 25.23 23.63 1.53 3.43 3.00 3.73 1.73 2.43 1.97 2.17 0.80 2.83
@@
Gemini-2.5-Flash-Thinking 92.71 86.87 8.67 9.27 8.70 9.00 7.80 8.93 8.90 9.00 8.03 8.57
@@
Claude-3.5-Sonnet 83.36 78.10 7.97 8.53 8.27 7.93 7.03 8.50 8.00 7.57 6.70 7.60
o1-mini-2024-09-12 78.96 73.93 7.27 8.43 7.90 7.53 6.27 8.27 8.17 6.43 6.63 7.03
GPT4 Turbo 78.57 73.60 6.97 8.17 8.33 7.80 6.00 7.57 8.13 7.00 6.43 7.20
gpt-4-0125-preview 76.44 71.60 6.90 7.40 8.03 7.30 6.00 7.47 7.63 6.87 6.33 7.67
Baidu4.0 75.08 70.33 7.27 7.23 7.67 7.43 5.63 6.47 6.80 7.63 7.80 6.40
Baidu3.5 69.09 64.73 6.20 6.70 7.80 6.83 5.20 5.50 6.00 7.23 6.57 6.70
ChatGLM-pro 69.09 64.73 5.90 7.07 7.03 7.90 5.43 6.33 5.00 6.67 5.97 7.43
gpt4-0613 66.16 61.97 6.50 6.73 6.60 6.73 5.43 6.10 6.47 5.30 5.20 6.90
Spark-v3.0 65.62 61.47 5.77 6.50 7.27 7.30 5.70 5.90 5.03 6.50 5.23 6.27
Claude-3-Haiku 62.97 58.97 5.80 6.60 6.97 6.63 4.83 5.93 6.33 4.80 5.23 5.83
Gemini Pro 58.18 54.50 4.87 5.43 7.07 6.43 5.10 4.50 4.65 6.33 4.42 5.70
GPT-3.5-turbo 55.41 51.90 4.97 5.37 6.40 6.47 4.43 4.67 5.43 4.20 4.37 5.60
@@
Model ZS FS CoT Mean Variance
Open-source LLMs
DeepSeek-R1 91.23 89.33 88.43 89.67 1.36
DeepSeek-V3 90.37 86.60 88.40 88.46 2.37
Qwen-3-235B 90.33 84.97 87.27 87.52 4.83
Qwen-3-32B 86.37 80.27 85.37 84.00 7.14
Closed-source LLMs
Doubao-1.5-Thinking-Pro 93.67 90.63 91.73 92.01 1.57
Gemini-2.5-Pro 91.07 87.57 90.60 89.74 2.41
Gemini-2.5-Pro-Thinking 91.00 87.87 90.23 89.70 1.78
Doubao-1.5-Pro 89.63 86.43 87.30 87.79 1.83
Claude-Sonnet-4-Thinking 85.27 85.60 86.87 85.91 0.48
Claude-Sonnet-4 85.23 84.77 86.23 85.41 0.37
GPT-4o 82.50 80.13 83.30 81.98 1.81
o1 87.43 86.07 86.87 86.79 0.31
o3-mini 78.80 76.97 79.67 78.48 1.27
Table 8: Model Capability under three prompting paradigms. ZS: Zero-Shot; FS: Few-Shot; CoT: Chain of Thought.
Appendix D Details in Experiment
@@ D.2 LLM-as-Judge Validation
Appendix E Implementation Details
This section provides detailed information about the annotation processes and evaluation procedures underlying our LLMEval-3 platform. We describe the expert involvement in data curation, validation processes, and the associated costs to ensure transparency and reproducibility.
E.1 Data Annotation Process
@@
E.2 Manual Evaluation Process
@@
E.3 Cost Analysis
The development and validation of LLMEval-3 required substantial investment in both human expertise and computational resources. We spent more than \$5,000 on using latest APIs of LLMs and deploying models for evaluation purposes. Additionally, \$10,000 was allocated for hiring qualified volunteers to conduct manual evaluations, ensuring rigorous validation of our automated assessment framework.
@@ E.3 Cost Analysis
Appendix F JWT Authentication Process
@@
Our JWT implementation follows a standard three-phase protocol: token generation, transmission, and verification. Algorithm 1 outlines the complete JWT authentication workflow used in LLMEval-3.
Algorithm 1 JWT Authentication Process in LLMEval-3

### v5

@@ Appendix E LLMEval-Fair Leaderboard
Qwen3-235B 96.42 90.32 9.23 9.43 9.03 9.50 8.23 9.43 8.97 9.17 8.73 8.60
QwQ-32B 94.51 88.53 8.30 9.46 9.23 9.33 7.83 9.46 8.65 9.27 8.57 8.43
DeepSeek-V3.2 92.27 86.43 8.73 9.13 8.53 8.70 7.40 9.33 8.87 9.37 8.53 7.83
Qwen3-32B 92.22 86.38 8.43 9.10 8.57 9.10 7.77 9.47 8.67 9.30 7.70 8.27
GLM-4-32B 88.43 82.83 7.77 8.97 8.33 8.33 7.03 9.13 8.27 8.77 8.23 8.00
Qwen2.5-32B-Instruct 85.06 79.68 7.70 8.57 8.33 8.33 6.70 8.50 8.17 7.70 7.60 8.08
Qwen-Turbo-1101 83.72 78.42 7.97 8.37 8.03 8.23 6.40 8.50 8.10 7.50 7.27 8.05
Yi-34B-Chat 70.15 65.71 5.77 6.63 7.37 7.53 5.47 5.77 5.47 7.47 6.30 7.93
Megrez-3B-Instruct 67.01 62.77 5.80 6.77 6.80 7.13 5.40 6.87 5.70 6.53 5.70 6.07
Qwen2-7B-Instruct 65.15 61.03 5.47 6.73 6.33 7.60 5.13 6.17 6.17 5.73 5.33 6.37
Nanbeige-Plus 65.10 60.98 5.78 5.57 6.77 7.37 5.37 5.93 5.45 6.30 5.67 6.77
Phi-4-Final 63.98 59.93 5.80 6.47 6.23 6.53 5.53 6.30 6.27 5.50 5.43 5.87
Llama-3.2-90B-Vision 61.74 57.83 5.63 6.33 6.20 5.80 4.73 6.10 6.57 5.03 5.27 6.17
Llama-3.3-70B 60.85 57.00 5.80 6.90 5.63 5.70 5.47 5.70 6.30 4.70 4.87 5.93
Baichuan2-13B-Chat 58.28 54.59 4.47 5.53 7.40 6.90 4.63 4.80 4.33 6.23 4.60 5.70
Qwen-plus 56.58 53.00 4.40 5.10 6.53 6.53 5.00 4.77 4.87 5.17 5.13 5.50
@@
Nanbeige-16B 55.45 51.94 4.37 5.30 6.50 6.30 3.97 4.70 4.07 5.90 4.73 6.10
Mixtral-8x7B-Instruct 51.69 48.42 4.27 5.47 6.47 6.40 3.13 4.50 5.07 3.57 4.37 5.17
ChatGLM-2-6B 42.31 39.63 2.33 3.77 5.97 6.13 2.83 3.83 2.60 3.80 4.00 4.37
Llama-3.1-8B 41.25 38.64 3.87 4.20 4.27 4.17 3.50 3.83 4.30 3.17 3.20 4.13
Ziya-13B-v1.1 40.18 37.64 2.77 3.97 5.17 5.33 2.80 3.77 2.53 3.70 3.03 4.57
InternLM-7B-Chat 38.71 36.26 2.63 3.67 4.87 5.57 3.17 3.33 2.33 4.03 3.13 3.53
Linly-LLaMA2-13B 37.03 34.69 2.20 3.77 4.50 5.00 2.43 3.33 2.53 3.90 2.50 4.53
Phi-3-Medium-128K 36.95 34.61 2.27 4.17 3.70 4.23 2.87 4.50 3.57 3.20 2.27 3.83
BELLE-Llama2-13B-Chat 36.25 33.96 2.57 3.07 4.93 4.73 2.83 3.80 2.43 3.33 2.40 3.87
Llama-2-7B-Chat 25.22 23.62 1.53 3.43 3.00 3.73 1.73 2.43 1.97 2.17 0.80 2.83
@@
Claude-Sonnet-4.5 93.31 87.40 8.80 8.97 8.93 8.73 8.37 9.10 8.97 8.93 8.13 8.47
Gemini-2.5-Flash-Thinking 92.74 86.87 8.67 9.27 8.70 9.00 7.80 8.93 8.90 9.00 8.03 8.57
@@
Claude-3.5-Sonnet 83.38 78.10 7.97 8.53 8.27 7.93 7.03 8.50 8.00 7.57 6.70 7.60
o1-mini 78.93 73.93 7.27 8.43 7.90 7.53 6.27 8.27 8.17 6.43 6.63 7.03
GPT-4-Turbo 78.57 73.60 6.97 8.17 8.33 7.80 6.00 7.57 8.13 7.00 6.43 7.20
GPT-4-Preview 76.44 71.60 6.90 7.40 8.03 7.30 6.00 7.47 7.63 6.87 6.33 7.67
Baidu-4.0 75.08 70.33 7.27 7.23 7.67 7.43 5.63 6.47 6.80 7.63 7.80 6.40
Baidu-3.5 69.10 64.73 6.20 6.70 7.80 6.83 5.20 5.50 6.00 7.23 6.57 6.70
ChatGLM-Pro 69.10 64.73 5.90 7.07 7.03 7.90 5.43 6.33 5.00 6.67 5.97 7.43
GPT-4-Legacy 66.15 61.96 6.50 6.73 6.60 6.73 5.43 6.10 6.47 5.30 5.20 6.90
Spark-3.0 65.62 61.47 5.77 6.50 7.27 7.30 5.70 5.90 5.03 6.50 5.23 6.27
Claude-3-Haiku 62.93 58.95 5.80 6.60 6.97 6.63 4.83 5.93 6.33 4.80 5.23 5.83
Gemini-Pro 58.18 54.50 4.87 5.43 7.07 6.43 5.10 4.50 4.65 6.33 4.42 5.70
GPT-3.5-turbo 55.42 51.91 4.97 5.37 6.40 6.47 4.43 4.67 5.43 4.20 4.37 5.60
@@
Appendix F Details in Experiment
@@ F.2 LLM-as-Judge Validation
Appendix G Implementation Details
This section provides detailed information about the annotation processes and evaluation procedures underlying our LLMEval-Fair platform. We describe the expert involvement in data curation, validation processes, and the associated costs to ensure transparency and reproducibility.
G.1 Data Annotation Process
@@
G.2 Manual Evaluation Process
@@
G.3 Cost Analysis
The development and validation of LLMEval-Fair required substantial investment in both human expertise and computational resources. We spent more than \$5,000 on using latest APIs of LLMs and deploying models for evaluation purposes. Additionally, \$10,000 was allocated for hiring qualified volunteers to conduct manual evaluations, ensuring rigorous validation of our automated assessment framework.
@@ G.3 Cost Analysis
Appendix H JWT Authentication Process
@@
Our JWT implementation follows a standard three-phase protocol: token generation, transmission, and verification. Algorithm 1 outlines the complete JWT authentication workflow used in LLMEval-Fair.
Algorithm 1 JWT Authentication Process in LLMEval-Fair